OpenStudy (anonymous):

The polynomial (x-a)^3+b is zero at x=1 and when divided by x the remainder is -7. Find the values of a and b.

OpenStudy (anonymous):

?=-7

OpenStudy (raden):

are you sure that's when divided by " x " ?

OpenStudy (anonymous):

what is the the number that divided by x to get -7

OpenStudy (anonymous):

there's no number, its just divided by x to get -7

OpenStudy (anonymous):

you mean (x-3)^3+4/x = -7 ???

OpenStudy (anonymous):

dean12

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

no no you cant get -7

OpenStudy (anonymous):

so it's not possible to solve?

OpenStudy (anonymous):

no the problem is on the div try to make sure of this Ex

OpenStudy (raden):

i have 2 equations according information above : (1 - a)^3 + b = 0 ... (1) (x - a)^3 + b = -7 ... (2) there are 3 variables a,b, and x, actually we need 3 equations to solve them

OpenStudy (anonymous):

how do we get a third equation?

OpenStudy (raden):

maybe subtract the 1st to 2nd equa, then cancel the b's... but not sure it works

OpenStudy (anonymous):

(x - a)^3 + b = -7 i cant understand this

OpenStudy (raden):

that's remainder thorem, @AntarAzri

OpenStudy (anonymous):

@RadEn what do you mean f(x) = -7 for example ?

ganeshie8 (ganeshie8):

when u solve it, u wil get the value of x at which f(x) leaves a remainder of -7

ganeshie8 (ganeshie8):

if that convinces u.. :)

OpenStudy (anonymous):

@ganeshie8 can you type it on the equation tool below

ganeshie8 (ganeshie8):

type what

OpenStudy (anonymous):

what you are saying use the equation tool

OpenStudy (raden):

i sure there is a missing number after word " when divided by x " ... that should says " when divided by (x + or - a) " , with a constant

OpenStudy (anonymous):

there isn't

ganeshie8 (ganeshie8):

maybe take a screenshot and post the original question.. that may help

OpenStudy (anonymous):

I've attached the question

ganeshie8 (ganeshie8):

there will be multiple values i think

OpenStudy (anonymous):

multiple values are possible most likely, but do you know of any way to solve it

ganeshie8 (ganeshie8):

cool, then we can try... let me think a bit i was thinking u looking for an uniq value of a and b

ganeshie8 (ganeshie8):

The polynomial (x-a)^3+b is zero at x=1 (1-a)^3+b = 0 ----------(1)

ganeshie8 (ganeshie8):

when divided by x the remainder is -7 when divided by x-0 the remainder is -7 that means f(0) = -7 (0-a)^3+b = 0 ---------(2)

ganeshie8 (ganeshie8):

2 equations, 2 unknowns u can solve them

ganeshie8 (ganeshie8):

from (2), b = -a^3 put this value in (1) (1-a)^3 - a^3 = 0 u wil get a quadratic in a, solve it

OpenStudy (anonymous):

I got the answer, thanks for your help!

ganeshie8 (ganeshie8):

np :) you must have gotten 2 value for a since its a quadratic